OpenStudy (anonymous):

Match the function (f) with its inverse (f^-1). f(x)=3x + 3

OpenStudy (anonymous):

here are the answer choices 3x - 12 f(x)-1 = x-3/3 3x + 12 This function does not have an inverse

OpenStudy (anonymous):

y=f(x)=3x + 3 sfr x then swap y and x

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

y=3x+3 y-3=3x subtract 3 y/3 -1=x div by x swap x and y y=x/3 - 1 that is your inverse

OpenStudy (anonymous):

so its f(x)-1 = x-3/3 ?

OpenStudy (anonymous):

(x-3)/3 make sure to use parenthesis

OpenStudy (anonymous):

yes sir!

OpenStudy (anonymous):

wat about \[f(x)=1/3x+4\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):